OpenStudy (anonymous):

(1-sin^2 x)/(1-cos^2 x) = cot^2 x work only on one side

OpenStudy (anonymous):

\[RHS=\frac{1-\sin^2x}{1-\cos^2x}\] \[=\frac{\cos^2x}{\sin^2x}\] \[\cot^2x=RHS\]

OpenStudy (anonymous):

*LHS at first

OpenStudy (anonymous):

got it, thanks!